# UAD: Unsupervised Affordance Distillation for Generalization in Robotic Manipulation

### v1

@@ Abstract
Understanding fine-grained object affordances is imperative for robots to manipulate objects in unstructured environments given open-ended task instructions. However, existing methods of visual affordance predictions often rely on manually annotated data or conditions only on a predefined set of tasks. We introduce Unsupervised Affordance Distillation (UAD), a method for distilling affordance knowledge from foundation models into a task-conditioned affordance model without any manual annotations. By leveraging the complementary strengths of large vision models and vision-language models, UAD automatically annotates a large-scale dataset with detailed $<$instruction, visual affordance$>$ pairs. Training only a lightweight task-conditioned decoder atop frozen features, UAD exhibits notable generalization to in-the-wild robotic scenes and to various human activities, despite only being trained on rendered objects in simulation. Using affordance provided by UAD as the observation space, we show an imitation learning policy that demonstrates promising generalization to unseen object instances, object categories, and even variations in task instructions after training on as few as 10 demonstrations. Project website: unsup-affordance.github.io/.
@@ IV-A Task-Conditioned Affordance Prediction
Methods KLD (↓) SIM (↑) NSS (↑)
Cross-View-AG [59] 1.787 0.285 0.829
LOCATE [124] 1.405 0.372 1.157
3DOI [125] 3.565 0.227 0.657
AffordanceLLM [126] 1.463 0.377 1.070
UAD (Ours) 0.526 0.366 1.359
TABLE I: Evaluation results on AGD20K test split.
For evaluation, since UAD conditions on free-form language and AGD20K contains only a list of pre-defined actions, we format the instructions as “region to <action> the <object>”. We follow the evaluation procedure defined in AGD20K [59] and evaluate UAD on the test split using the same metrics as in previous work, namely KL Divergence (KLD), Similarity Metric (SIM), and Normalized Scanpath Saliency (NSS). Most of these metrics consider affordance prediction as an image-level distribution rather than a pixel-level distribution. As a result, we adopt a simple normalization and observe that while UAD is not trained on a similar distribution, UAD still performs competitively compared to prior work. Interestingly, UAD can even generalize to a number of human activities that involve objects and task instructions completely out of distribution from our training set, such as “eating bananas”, “taking photos”, “sitting on bicycles”, “holding golf clubs”, “lying on bed”, and “typing on computers”, as shown in Fig. 5.
@@ A-A Details on Task-conditioned Affordance Prediction (Sec IV-A)
Details for Evaluation on AGD20K Dataset. We evaluate our model on the easy split of AGD20K and compare with the baseline performance reported in [126]. To avoid numerical instability in KLD computation as we consider per-pixel affordance instead of per-image affordance as in other works, we post-process our model’s prediction by adding a small $\epsilon$ to each pixel before normalization.
@@ A-B Policy Learning in Simulation Details (Sec IV-B)
We use a key-frame based policy for both demonstrations and learnt policies. Key-frames are commonly used by prior works [128, 121, 122, 129, 130] as “important or bottleneck steps of gripper during task execution”. To execute an action $<$end-effector pose, gripper action$>$, we first command the end-effector controller of an interpolated trajectory from current to target pose, then execute the gripper action afterwards.
@@
During training, we normalized the channels for visual observation by: normalize visual representation to $(-1,1)$, clip $(x,y,z)$ to the min and max workspace bounds, and set depth for all out-of-bound points to 0. Additionally, we append channels according to pixel location, following the original implementations in RVT. We apply random cropping augmentation to visual input during each training step.
@@ A-C Policy Learning in Real-World Details (Sec IV-C)
Environment Setup. Our real-world evaluation platform uses a Franka arm mounted in a tabletop setup built with Vention frames. Since the learned policy outputs 6-DoF end-effector poses and gripper actions, we use position control in all experiments, which is running at a fixed frequency of $20$ Hz. Specifically, given a target end-effector pose in the world frame, we first clip the pose to the pre-defined workspace. Then we perform linear interpolation from the current pose of the robot to the target pose with a step size of $5$mm for position and $1$ degree for rotation. To move to each interpolated pose, we first calculate inverse kinematics to obtain the target joint positions based on current joint positions using the IK solver implemented in PyBullet [131]. Then we use the joint impedance controller from Deoxys [132] to reach to the target joint positions. Two RGB-D cameras, Orbbec Femto Bolt, are mounted on the left side and the right side of the robot facing the workspace center. The cameras capture RGB images and point clouds at a fixed frequency of $20$ Hz.

### v2

@@ Abstract
Understanding fine-grained object affordances is imperative for robots to manipulate objects in unstructured environments given open-ended task instructions. However, existing methods of visual affordance predictions often rely on manually annotated data or conditions only on a predefined set of tasks. We introduce Unsupervised Affordance Distillation (UAD), a method for distilling affordance knowledge from foundation models into a task-conditioned affordance model without any manual annotations. By leveraging the complementary strengths of large vision models and vision-language models, UAD automatically annotates a large-scale dataset with detailed <<instruction, visual affordance>> pairs. Training only a lightweight task-conditioned decoder atop frozen features, UAD exhibits notable generalization to in-the-wild robotic scenes and to various human activities, despite only being trained on rendered objects in simulation. Using affordance provided by UAD as the observation space, we show an imitation learning policy that demonstrates promising generalization to unseen object instances, object categories, and even variations in task instructions after training on as few as 10 demonstrations. Project website: unsup-affordance.github.io/.
@@ IV-A Task-Conditioned Affordance Prediction
Methods KLD (↓) SIM (↑) NSS (↑) NSS-0.5 (↑)
Cross-View-AG [59] 1.787 0.285 0.829 -
LOCATE [124] 1.405 0.372 1.157 1.723
3DOI [125] 3.565 0.227 0.657 -
AffordanceLLM [126] 1.463 0.377 1.070 -
UAD (Ours) 1.878 0.407 1.092 2.050
TABLE I: Evaluation results on AGD20K Unseen test split.
For evaluation, since UAD conditions on free-form language and AGD20K contains only a list of pre-defined actions, we format the instructions as “region to <action> the <object>”. We follow the evaluation procedure defined in AGD20K [59] and evaluate UAD on the Unseen test split using the same metrics as in previous work, namely KL Divergence (KLD), Similarity Metric (SIM), and Normalized Scanpath Saliency (NSS). We additionally report NSS-0.5, which uses a stricter threshold (0.5 vs. standard 0.1) to focus evaluation on the most salient ground truth regions. See Appendix  IV-A for details. Despite not being trained on AGD20K data, UAD performs competitively, achieving the highest SIM score and outperforming LOCATE on NSS-0.5. The higher KLD stems from UAD’s fine-grained predictions compared to AGD20K’s diffuse ground truth distributions. Interestingly, UAD can even generalize to a number of human activities that involve objects and task instructions completely out of distribution from our training set, such as “eating bananas”, “taking photos”, “sitting on bicycles”, “holding golf clubs”, “lying on bed”, and “typing on computers”, as shown in Fig. 5.
@@ A-A Details on Task-conditioned Affordance Prediction (Sec IV-A)
Details for Evaluation on AGD20K Dataset – Evaluation setting and Processing. We evaluate our model on the Unseen test split of AGD20K and compare with the baseline performance reported in [126]. To avoid numerical instability in KLD computation as we consider per-pixel affordance instead of per-image affordance as in other works, we post-process our model’s prediction by clipping per-pixel prediction to $\epsilon,1-\epsilon]$ before normalization.
We format instructions as ”region to action the object” to match AGD20K’s action vocabulary. However, AGD20K’s ground truth annotations contain ambiguities in some action-region mappings or select one specific region when multiple valid options exist. For instance:
  • For ”hit” with an axe, the ground truth highlights the handle (for gripping during hitting) rather than the head (which does the hitting).
  • For ”hold” with a cup, the ground truth exclusively marks the handle, though the cup body is also commonly used for stable holding.
Since UAD was not trained on AGD20K data, it lacks knowledge of these dataset-specific conventions. To ensure fair evaluation, we disambiguated these instructions to match the ground truth interpretations:
  • hit - axe → handle of axe to hold during hitting
  • ride - bicycle → region to sit on and push the bicycle
  • pour - cup → handle of the cup to hold while pouring
  • wash - cup → rim of the cup to wash
  • hold - cup → handle to hold the cup
Details for Evaluation on AGD20K Dataset – NSS-0.5 metric. The NSS metric evaluates model predictions at“fixation pixels”, defined as locations where ground truth affordance values exceed a specified threshold. It computes the average prediction at these pixels, with all pixels weighted equally. The standard evaluation uses a relatively low threshold (0.1). However, the AGD20K ground truth is generated by computing Gaussian mixtures centered on manually annotated keypoints, causing the affordance signal to diffuse beyond actual object boundaries into background regions. As visualized in Fig.8, the green areas represent pixels with GT >> 0.5, while the red areas show pixels with GT values between 0.1 and 0.5, which often extend into the background and may not correspond to actual object affordances. Under this original evaluation protocol (threshold=0.1), models are penalized for predicting low affordance in these peripheral regions (red), even though they may not correspond to actual object affordances. By using threshold=0.5, NSS-0.5 focuses evaluation on the core affordance regions (green) that are more strongly associated with the query. Our results show that UAD outperforms LOCATE—the previous state-of-the-art on standard NSS—on this more stringent metric, demonstrating our model’s ability to identify salient affordance regions.
Figure 8: Comparison of ground truth fixation regions at different NSS thresholds. Green regions indicate pixels with GT >0.5>0.5 (used for NSS-0.5), while red regions show pixels with GT values between 0.1 and 0.5 (additionally included in standard NSS).
Note on KLD scores. The KL Divergence metric heavily penalizes predictions that approach zero where ground truth values are positive. As shown in our threshold visualization, the AGD20K ground truth extends into background regions due to its Gaussian mixture construction. Since UAD’s predictions are focused on foreground objects and predict near-zero values in these peripheral areas where the ground truth retains small positive values, this partially contributes to UAD’s higher KLD score.
@@ A-B Policy Learning in Simulation Details (Sec IV-B)
We use a key-frame based policy for both demonstrations and learnt policies. Key-frames are commonly used by prior works [128, 121, 122, 129, 130] as “important or bottleneck steps of gripper during task execution”. To execute an action <<end-effector pose, gripper action>>, we first command the end-effector controller of an interpolated trajectory from current to target pose, then execute the gripper action afterwards.
@@
During training, we normalized the channels for visual observation by: normalize visual representation to (1,1)(-1,1), clip $(x,y,z)$ to the min and max workspace bounds, and set depth for all out-of-bound points to 0. Additionally, we append channels according to pixel location, following the original implementations in RVT. We apply random cropping augmentation to visual input during each training step.
@@ A-C Policy Learning in Real-World Details (Sec IV-C)
Environment Setup. Our real-world evaluation platform uses a Franka arm mounted in a tabletop setup built with Vention frames. Since the learned policy outputs 6-DoF end-effector poses and gripper actions, we use position control in all experiments, which is running at a fixed frequency of $20$ Hz. Specifically, given a target end-effector pose in the world frame, we first clip the pose to the pre-defined workspace. Then we perform linear interpolation from the current pose of the robot to the target pose with a step size of 55mm for position and $1$ degree for rotation. To move to each interpolated pose, we first calculate inverse kinematics to obtain the target joint positions based on current joint positions using the IK solver implemented in PyBullet [131]. Then we use the joint impedance controller from Deoxys [132] to reach to the target joint positions. Two RGB-D cameras, Orbbec Femto Bolt, are mounted on the left side and the right side of the robot facing the workspace center. The cameras capture RGB images and point clouds at a fixed frequency of $20$ Hz.